# ObjectFinder: Open-Vocabulary Assistive System for Interactive Object Search by Blind People

### v1

@@ 4.1.1. Participants and Procedure
The interview was conducted at the beginning of the exploratory user study described in Sec. 5. The participants were compensated €$40$ for study participation, and travel costs were reimbursed.
4.1.2. Data Analysis
We analyzed qualitative data using the thematic analysis framework developed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006). The first author led the analysis, first reading and re-reading the transcripts, then engaging in coding, and finally organizing the data into themes. The themes were subsequently reviewed by the research team. In this phase, $122$ data points were categorized into $42$ initial codes, which were then distilled and refined into $15$ codes, from which two main themes were crafted.
4.2. Results
Here, we present the two themes crafted from the Phase II interviews, (1) Challenges of Existing Technology for Object Search, and (2) Target Objects and Difficult Cases for Object Search.
4.2.1. Challenges of Existing Technology for Object Search.
The data in Table 1 shows that most participants prefer description-based applications. Notably, only participant P4 reported using the detection-based application Lookout. The use of Be My Eyes for document summarization and volunteer interaction is not included in this table.
Description-based Systems
One challenge for blind people is to take the right photo for a description-based system due to the heavy reliance of photography on the visual domain (C1). One participant commented on Be My Eyes, “if you lose this point, you can receive the wrong truth. A wrong picture you never can be sure, you have to know it if you use these things. ” (P2). Seeing AI and Envision do not help to take the right photo as well, but they do indicate if the photo is of low quality. “There is a Seeing AI app for example. It tells you, ‘No, this corner is not visible’. […] Envision, it doesn’t tell you unless you already take the photo and it processes the photo, then it tells you ‘no it was quite blurry’.” (P5). In terms of learning if the users of Be My Eyes (P0-3) can align the photo with the real environment: all of them answered yes, but participants reported that it is difficult in an unknown environment, e.g., “it’s easier when I know the surroundings, so I can align what’s described. But somewhere I’ve never been before, I find it more difficult.”(P1). Besides, P2 mentioned a practical method to take good photos: “ you can learn to fix: you can go [to] touch the object, and you go one or two steps back, take your camera and focus.”. This suggests that the effectiveness of description-based systems in developing a mental map within an unfamiliar environment without tactile exploration may be limited. In addition, participants were suspicious of AI-generated descriptions (Lee et al., 2022a; Xiao et al., 2024) without knowing “when does it start lying?” (P2). However, the availability of these applications on the cell phone was considered a significant advantage (P1, P3), while privacy concerns must be considered, as everything is “sent to the cloud” (P0).
Other methods for Object Search
As for object-search systems, P2 noted: “at the moment we have no usable object-finding system in most cases.”. However, participants mentioned various methods to search for objects. For example, P4 uses Lookout, but not frequently due to poor internet connection. Some mentioned using tactile exploration and white canes to search for objects in their daily activities (P2, P3, P7). As for other technologies, P2 mentioned adding audio descriptions to QR codes with PenFriend to distinguish private items, such as clothing colors, food containers, or medication. Besides, participants attached AirTags to their keys, although not everything can be equipped with the technology (P2, P6). It enables participants to locate items using sounds. For example, they might use an Apple Watch to ping their phone to find their seat in a beer garden. However, the aforementioned methods do not have the flexibility to define targets flexibly, and cannot inspire users with interesting unexpected discoveries in unfamiliar scenarios (C2).
4.2.2. Target Objects and Difficult Cases for Object Search.
Here, we analyse the role of target objects and difficult cases for object search in both familiar and unfamiliar scenarios.
Familiar Scenarios
In their familiar private spaces, the belongings of participants are well-managed and placed in specific locations. However, search for them can be difficult if objects are moved by other people (P2, P5), or if participants misplace their own items (P5, P8). As P2 mentioned, “ someone was in my office and changed a little thing to put her laptop and something on the table. After that, it took an hour to have my things back.”. P8 also noted, “ if I’m lazy and put it somewhere, then I really know it has to be around here, but it’s not in its usual place.” Participants also reported searching for small objects in their daily lives: More than one participant had previously looked for their smartphone, keys, small electronics (charger, USB stick, etc.), or items such as coffee cups. Sometimes, even though the participants know that objects are nearby, exact location matters, with one participant pointing out that “‘Near’ could be a problem […] your coffee spoiled your laptop.” (P2)
Unfamiliar Scenarios
In unfamiliar environments, participants reported difficulties in knowing what is on top of a table (P3, P5, P8) , such as food, because “you usually don’t feel around whole tabletop.” (P3). In addition, the participants are unlikely to know the paintings and displays hanging on the wall (P0) and the “hidden” objects (P6).
Additionally, knowing what one can use in an unknown scenario can be a challenge (C2). For example, P4 explained that “we once asked ourselves in the student room whether we had air conditioning. […] The question really occupied us for several hours [and] whether it would be warm in the practice room in the summer. […] If we could scan the room ourselves, we could determine ourselves whether climate regulation is present.” In contrast to the typically small target objects in their daily lives, most participants reported looking for large pieces of furniture when they did not have a mental map of an unknown environment , because “I find it good to get oriented quickly, especially in a foreign room.” (P6). P1 typically aimed to find the “next available chair” in an unfamiliar office. Meanwhile, P2 sought to locate the cupboard, bed, and other furniture in an unfamiliar hotel room, noting that plugs were harder to identify. In smaller rooms, participants commented on the value of tactile exploration to investigate their surroundings, with P7 noting that he “go through the room and feel with my hands to see what I can find next.” However, it is important that this is not feasible in all environments. For example, in the specific case of an unfamiliar toilet, “it’s very important to find the objects, because [otherwise] you have to touch everything” (P2).
Functions Passive Information Aquisition Interactive Object Search
Lookout (Google, 2024) Be My Eyes (Eyes, 2024) ObjectFinder_Base ObjectFinder
@@ 4.2. Results
[…]
Figure 6. Simplified user study procedure focusing on two target objects: a large piece of furniture (an office chair) and a smaller object (a plate of cookies). The user walks into an unknown environment, the office in this example. To sit in front of a desk, he or she should find an office chair first and navigate to it. Then the user sits on the office chair and defines the desk as the region of interest. The user will know what is on the desk through the scene description function, a plate of cookies in this example, and get additional information through open questions.
[…]
5. Phase III: Exploring User Perspectives on ObjectFinder
In Phase III, we conducted an exploratory study with blind people to determine if ObjectFinder can support the object search in unfamiliar environments. In this step of our work, we include our own prototype ObjectFinder, ObjectFinder_Base (a closed-vocabulary baseline prototype), and the commercial systems Be My Eyes (Eyes, 2024) (a description-based application) and Lookout (Google, 2024) (a detection-based application) as points of reference for participant feedback.
5.1. Method
We invited participants P1111-P8888 (Sec. 5.1.1) to explore the two prototypes and the commercial applications in two unknown environments that were set up in our research lab: a space mimicking a living room, and an office. The functional differences between the tested systems are listed in Table 2. ObjectFinder_Base differs from ObjectFinder that it only recognizes a limited number of objects with fixed wording. The systems are described more detailed in Sec. 5.1.2.
In order to assess users’ experiences with the systems and to find out whether an open-vocabulary offers an advantage over a closed-vocabulary approach, we compared ObjectFinder with ObjectFinder_Base in terms of: expectations regarding the functionality, mental map that users formed of the environment after using one system, cognitive load and satisfaction. We used semi-structured interviews featuring Likert-scales for function evaluation and open ended questions. Questions ranged from “In which areas do you believe our system could be improved?” and “Do you want more or less information?”. Additionally, we included the NASA-TLX (Hart, 2006) for assessing the cognitive load to accomplish the tasks, and the Net Promoter Score (NPS) (Reichheld, 2003), a measure that reflects the likelihood if the users would like to introduce the system to their community. The exploration of the commercial applications was conducted separately, followed by a short semi-structured interview and the Net Promoter Score (Reichheld, 2003).
5.1.1. Participants and Procedure
The participants were the same eight people who took part in the interview described in Phase II (see Sec.  4.1.1). For scene understanding, six of the participants had previous experience using description-based applications such as Seeing AI, Be My Eyes, and Envision, while only one participant uses a detection-based application, Lookout. In Table 1, we consider only the scene description or exploration feature of these applications, while usage of the applications for other purposes, such as reading documents, is not included.
[…]
Participants were compensated with €$40$ for study participation, and reimbursed for the travel cost.
5.1.2. Apparatus
In this section, we describe the systems we used in the evaluation: ObjectFinder, ObjectFinder_Base, Be My Eyes and Lookout.
ObjectFinder and ObjectFinder_Base. The hardware and flowchart of ObjectFinder and ObjectFinder_Base are identical, except that the baseline only detects object categories in COCO2017 (Lin et al., 2014). For ObjectFinder_Base, we set the confidence level threshold to $0.7$ to filter out falsely detected objects. In contrast, the YOLO-World used in ObjectFinder can detect any target outside the training dataset, resulting in potentially lower confidence levels. We set the confidence threshold for YOLO-World to $0.3$ and tested it with $20$ objects, including those used in our user studies. It detected all targets without any false positives, even when the objects were located $8$ meters away. Since some target objects are undetectable by the baseline system, we added additional options between blocks 1 and 2, as shown on the right side of Figure 5, to allow participants to trigger scene descriptions multiple times and determine if the target objects are present in the region in front of them.
Lookout. The Explore mode of Lookout, which was tested in this study, detects the nearest object in front of the camera and provides real-time outputs including object categories, quantities, and associated text.
Be My Eyes. Be My Eyes integrates GPT-4 into its Be My AI feature for a vivid scene description, replacing the need for sighted volunteers.
Lookout and Be My Eyes reflect the two main parts of our prototype: object detection and scene description. However, users cannot actively select their targets, obtain localization information and navigation instructions with these two applications, but instead passively receive information chosen by the system.
5.1.3. Scenario Exploration
@@
In each scenario (living room or office), participants were asked to find six target objects: three large pieces of furniture, in order to build a mental map of the unfamiliar environment, followed by three smaller objects found on the desk or coffee table. Figure 6 describes a simplified procedure involving the five functions F1-F5. Table 3 specifies the target objects that need to be found in sequence and the initial MLLM functions to be triggered when these target objects are detected. The layout and the order of targets for search are shown in Figure 7(b). We are the first to engineer prompts that generate route planning instructions for blind individuals, guiding them to objects. Therefore, our primary focus is on testing the route planning function (F3). Since the closed-vocabulary ObjectFinder_Base can only detect a limited number of target objects (Lin et al., 2014), we categorize the six target objects in each scenario into three groups: unrelated to (two objects), related to (three objects), and exact in (one object) COCO2017. As for the related targets, we will hint to the participants to look for the related objects in COCO2017 when they are using the baseline, but they will experience a vocabulary gap between the target definition and MLLM information. As for the unrelated objects, the participant cannot even define the target objects (C1). So we provide the option that the participant can trigger the scene description function several times to find the objects, as illustrated on the right side of Figure 5. Cookies are the bonus target, and we observe whether the participants can recognize it themselves to validate the capability of our system to inspire the detection of unexpected targets (C2). Figure 8 illustrates several examples of how the participants detected the target objects and received the outputs of the system during the user study.
@@
The commercial applications, Lookout and Be My Eyes, which were already familiar to the participants and not specifically designed for object search tasks, were briefly used in the living room scenario to locate a fan and a teapot as a way to remind participants of their functions.
@@
Figure 8. Examples of functions F3-F5 include triggering route planning (F3) to reach both a large piece of furniture and a small item, using scene description (F4) to describe the coffee table, and employing open questions (F5) to inquire about details. Postures (first column) and key frames (second column) are visualized, with audio input and output also provided.
@@
5.1.4. Data Analysis
In this phase, we have both qualitative data and quantitative data. For qualitative data, the user study transcripts were analyzed using the hybrid process of inductive and deductive thematic analysis proposed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006). Analysis was led by the first author who read the transcript repeatedly to achieve familiarisation, followed by multiple rounds of coding. Beyond data-driven inductive coding, we also applied deductive coding where it yielded meaningful insights into the system’s capacity to identify regions of interest (C1) and to facilitate the discovery of unexpected targets (C2). In a workshop, the research team assigned $243$ data points to $69$ codes, which were further refined to $12$ codes, and finally, four themes were crafted and will be presented in Sec. 5.2.
Due to the small size of the user group, inferential statistical analysis of the quantitative data is not meaningful. Therefore, we limit our comparison to descriptive statistics.
5.2. Results
Here, we present the following four key themes crafted from the interview findings, which focus on the users’ Information Needs for Object Search, their experiences with Mental Mapping, their System Perception and Preferences as well as System Requirements.
5.2.1. Information Needs for Object Search
In this section, we analyse the amount of information the participants preferred, and the obligatory and optional information for object search.
@@ 5.2.1. Information Needs for Object Search
Regarding the amount of information provided, three participants found it adequate for their needs. However, when asked about their preference for more or less information, responses were divided: half preferred more, while the other half favored less. We noted that preferences for information quantity relate to individual processing styles: while some participants could ignore excess information, others felt overwhelmed. The participants chose to receive more information, generally expressing a preference for as much as possible. As one participant noted, “as much as you can get. It’s completely blank for me, so the more I have, the better.” (P5). P4 expressed that “things that don’t interest me, I can just ignore.” Conversely, P6 and P7 explained their preference for less information, attributing it to not being accustomed to processing such a large amount of visual information and feeling overwhelmed by it. P8 suggested that the information provided could perhaps be reduced after getting the first overview: “it could maybe be a little bit less, so if you know once there are some objects on the table, then you don’t really need this information the second or third time unless you ask the system what’s on the table.” Additionally, we observed that the amount of information varies across functions. Be My Eyes, used solely for scene description, received praise from three participants for its “detailed” information, although one deemed it excessive. In contrast, participants noted that our ObjectFinder delivered more information than necessary for efficient navigation.
@@
Participants (P6-8) highlighted the importance of localization information (distance and direction) after scene exploration. The interpretation of the distance should be clear and intuitive. P1 highlighted the usefulness of different distance measures noting, “I think the steps are good for like when it’s really near, so it’s just a few steps. But if there are longer distances. I think sometimes meters might be more useful […] So it sometimes has like $0.1$ meter […] you could have said ‘right in front of you’ or ‘one step’ like that.” Participants mentioned that the distances reported were inaccurate, appearing larger than they perceived. Upon reviewing the video, we observed that some discrepancies were caused by measurements taken from the head to the object, rather than horizontally, as shown in Figure 11. This inaccuracy may be more pronounced in our small indoor settings. To describe the direction, P5 suggested to “specify a $20$ centimeters margin”, noting that if an object is $10$ centimeters to the left or right, it’s still considered in the front.
Contextual information around the target is crucial for object search, such as relationships among surrounding objects and their location information. For example, P3 noted, “it’s good that the objects around it are announced, so you have some idea where the object you want to find is in relation to other objects.” A description should also include the distance which says something about the user as a reference point to the object. P8 suggested to improve ObjectFinder by incorporating distance into the MLLM output, “ when you make the description, it holds everything but it does not really talk about the distance.”
@@
Figure 10. Net Promoter Scores (NPS) for different systems, assessed by participants P1-P8 while exploring scenarios.
[…]
The bar chart of the net promoter scores for different systems: ObjectFinder (M=6.75, SD=2.66), ObjectFinder_Base (M= 5.88, SD=2.42), ObjectFinder with efficient hardware (M=8.13, SD=2.01), Be My Eyes (M=7.25, SD=3.01), Lookout (M=5.31, SD=2.43).
Figure 11. Special cases. Upper half: Initially, the participant did not detect the coffee table as an obstacle to the sofa. Subsequently, the participant stepped backward to detect the coffee table instead of lowering the head. Bottom Left: The participant needed cues to lower the head to detect items on the desk. Bottom Right: The participant felt uneasy about the distance ($1.5$ meters) to the office chair because the measurement was taken from the head to the chair, not horizontally.
[…]
This visualization depicts the postures and key frames for special cases where participants needed cues to lower their heads, and felt uncertain about the distance. In the example shown in the lower right part of the figure, a participant who is about 1.7 meters tall detects a chair in her vicinity (0.5 meters away), yet the distance is calculated to be 1.5 meters from the glasses to the chair.
5.2.2. Mental Mapping
There is some evidence that participants obtained more detailed environmental information through the use of ObjectFinder. P1 noted:“I think it was very good [to know] where I am. When it said where my target object was, it also told me the surrounding objects, because maybe if I find the surrounding object first, then I know, OK I’m close. ” P6 noted the benefits of discovering other usable items (C2) and understanding their arrangement, “I didn’t know there was another chair there. It gives more information about objects and their arrangement.” Participants wished to get an overview of unfamiliar rooms without receiving too many details. P8 explained, “I could imagine if you really don’t know the room, and you want to first [have an] overview what is in the room, it was pretty detailed. ”
@@ 5.2.2. Mental Mapping
This highlights the potential of ObjectFinder to support users in building a mental map of an unknown space. This was more difficult using the existing commercial systems. For example, after using Be My AI, P8 mentioned, ”Be My AI had no information regarding the distances, so you have […] a less sense of location ”. As previously analyzed, localization information (distance and direction) and the relationships between surrounding objects are crucial for effective object search. The lack of localization information (P2, P8) and the absence of context regarding object relationships (P2) were mentioned as reasons why participants did not prefer Lookout, highlighting the potential for solutions that can combine multiple aspects of object search.
5.2.3. System Perception and Preferences
In this section, we analyze the general perception of ObjectFinder, the user preference with the systems, and the cognitive load to finish the tasks.
@@ 5.2.3. System Perception and Preferences
Generally, five out of eight participants expressed excitement about ObjectFinder after exploring it. “I was very positively surprised at how good it works, and how easy it is to get information and to find myself in the scene.” (P1). ObjectFinder was also appreciated for accurate object searches without physical touch. Subsequently, participants rated the system’s helpfulness ($\mu=4.13,\sigma=0.641$) and their independence ($\mu=4.31,\sigma=1.01$) with ObjectFinder in an unfamiliar environment, with both ratings based on $5$-point Likert scales and averaged over Through ObjectFinder, participants discovered more than just target items, finding inspiration in unexpected objects, both small and large (C2). For example, “I’m sort of inspired because now I know what options I have, instead of walking in and just looking for my phone because I know it must be there, but I don’t know what else could be there […] I was inspired to look for, for example, bananas and stuff like that.”. (P1). However, P3 noted that searching with ObjectFinder might be slower than tactile exploration in our compact indoor setting, noted, “there are so many steps putting it up and searching for the thing, and maybe it doesn’t even find it (ObjectFinder_Base).” The search could be accelerated by using ObjectFinder alongside the cane, taking advantage of the cane’s large radius (P2).
@@
The NPS for the prototypes does not vary substantially between the scenarios, office ($\mu=6.25,\sigma=2.49$) and living room ($\mu=6.38,\sigma=2.67$), with scores being generally high, and not lagging behind those allocated to commercially available systems (see Figure 10). With respect to differences between the systems and implications for user preferences, participants appreciated that they could actively define the target objects with ObjectFinder, making it more reliable in searching for specific items. As P6 mentioned, “what I also liked about this system is speaking to it, I find that more targeted.”. P3 echoed this sentiment, highlighting the biggest advantage: “The biggest advantage […] contrary to Be My AI is that you can specify what you want to find.” (C1).
As shown in Table 4, the study also examined cognitive load, showing that it was relatively low across system types and scenarios. Retrieving keyframes for MLLM information and participant feedback reveals that the landscape orientation of the camera on the glasses, contrasting with participants’ usual practice of taking portrait photos using commercial applications, and the presence of relatively low furniture in the living room often result in the prototype failing to recognize the coffee table as an obstacle to the sofa, as illustrated in Figure 11. The exploratory experience with the prototype was not comprehensive enough. P7 mentioned that F4 and F5 are insufficiently tested in our study, and added: “one needs to get used to it, it requires some practice, and then the search will be very precise.” P5 mentioned that she was not accustomed to the earcon indicating to stand still.
@@
5.2.4. System Requirements
In this section, we analyze the system requirements, i.e. interaction features, software and hardware requirements.
@@ 5.2.4. System Requirements
Additionally, P8 described the voice commands as “pretty, pretty easy”, while P4 expressed concerns with programs that solely accept voice input, suggesting “maybe it’s better to have the ability to switch between a list of maybe recognized objects and voice commands.” Besides the current earcons, P6 suggested creating an additional earcon for option confirmation, rather than using spoken text.
@@
We asked participants to rate the functions before and after using ObjectFinder and ObjectFinder_Base in the explored scenarios, in terms of their perceived importance and level of interest, as captured by Likert scores (Figure 9). From the ratings, the functions object detection, localization information and route planning are important for the object search task ($\mu{>}4.0$), accompanied by a reduced standard error. Among the five functions evaluated, localization information was rated as the most important.
Compared with the commercial applications, P8 mentioned the advantage of ObjectFinder is “the distance and also the guiding function which is not really available for the other both apps.” (P8). We note that the software should be fast and make fewer mistakes. P2 mentioned that receiving explanations from a sighted person on Be My Eyes “works better because it’s without delay.” Half of the participants mentioned frequent misidentifications of Lookout was not preferred. As P5 noted, “it misidentified objects […], like there’s no dishwasher,” and at times, the teapot was categorized as a helmet or mouse.
@@
Five participants mentioned that the hardware was heavy for long-term use. We asked participants to imagine that the ObjectFinder prototype was implemented on more efficient hardware (Waisberg et al., 2024), such as the Ray-Ban glasses for image capturing paired with an iPhone for data processing, while actually trying on a pair of Ray-Ban glasses to evaluate their feel. With these enhancements, the Net Promoter Score (NPS) of ObjectFinder increased from ${\mu}{=}6.750,{\sigma}{=}2.659$ to ${\mu}{=}8.125,{\sigma}{=}2.013$, as illustrated in Figure 10. In comparison, a common advantage of the applications, Be My AI and Lookout, is the availability on the smartphone.

### v2

@@ 3.4. Function Implementation
4. User Evaluation
In order to understand how ObjectFinder can support object search in unfamiliar environments, we conducted an exploratory study with eight blind people. In this step of our work, we include our own prototype ObjectFinder, ObjectFinder_Base (a closed-vocabulary baseline prototype), and the commercial systems BeMyAI (Eyes, 2023) (a description-based application) and Lookout (Google, 2024) (a detection-based application) as points of reference for participant feedback. Specifically, we focus on the following research questions:
  1. RQ1:
    To what extent does ObjectFinder deliver the necessary information for effective object search?
  2. RQ2:
    How do blind people perceive the detailed scene context that ObjectFinder provides to facilitate object searches?
  3. RQ3:
    How do blind individuals perceive ObjectFinder generally and in comparison to description- and detection-based systems?
  4. RQ4:
    What requirements do blind users consider important for an object search system, as their experiences with ObjectFinder show?
4.1. Participants and Procedure
@@
4.2. Scenario Exploration
@@
Blind people typically search for large items as landmarks to construct mental maps of unfamiliar environments, and they would like to use the system to explore small objects on the tabletop. In each scenario (living room or office), participants were asked to find six target objects: three large pieces of furniture to establish spatial understanding, followed by three smaller objects found on the coffee table or desk. Figure 7 describes a simplified procedure involving the five functions F1-F5. Table 3 specifies the target objects that need to be found in sequence and the initial MLLM functions to be triggered when these target objects are detected. The layout and the order of targets for search are shown in Figure 8(b). We are the first to engineer prompts that generate route planning instructions for blind individuals, guiding them to objects. Therefore, our primary focus is on testing the route planning function (F3). Since the closed-vocabulary ObjectFinder_Base can only detect a limited number of target objects (Lin et al., 2014), we categorize the six target objects in each scenario into three groups: unrelated to (two objects), related to (three objects), and exact in (one object) COCO2017. As for the related targets, we will hint to the participants to look for the related objects in COCO2017 when they are using the baseline, but they will experience a vocabulary gap between the target specification and MLLM feedback. As for the unrelated objects, the participants cannot even specify the target objects (C1). So we provide the option for the participants to trigger the scene description function several times to find the objects. Cookies are the bonus target, and we observe whether the participants can recognize it themselves to validate the capability of our system to inspire the detection of unexpected targets (C2). Figure 9 illustrates examples of how the participants detected the target objects and received the system feedback during the user study.
@@
Lookout and BeMyAI, commercial applications familiar to participants but not designed for object search, were briefly used in the living room scenario to locate a fan and a teapot as reminders of their functions.
@@
Figure 9. The first column shows examples of postures, while the second column displays key frames captured when target objects were detected (F1) along with the corresponding real-time egocentric localization information (F2). The third column presents system feedback generated by the MLLM, which uses route planning (F3) to reach both large and small items, employs scene description (F4) to describe the coffee table, and utilizes open questions (F5) to gather additional details.
@@
4.3. Data Analysis
We have both qualitative and quantitative data. For qualitative data, the user study transcripts were analyzed using the hybrid process of inductive and deductive thematic analysis proposed by Fereday and Muir-Cochrane (Fereday and Muir-Cochrane, 2006). The first author led the analysis by repeatedly reading the transcript for familiarization and coding it in multiple rounds. Beyond data-driven inductive coding, we also applied deductive coding, which yielded meaningful insights into the system’s capacity to identify regions of interest (C1) and to facilitate the discovery of unexpected targets (C2). In a workshop, the research team assigned $243$ data points to $69$ codes, which were further refined to $12$ codes, and finally, four themes were crafted and will be presented in Sec. 5. For quantitative data, we limit our comparison to descriptive statistics due to the small size of the user group.
5. Findings
5.1. RQ1: To what extent does ObjectFinder deliver the necessary information for effective object search?
Participants found that ObjectFinder provides an adequate amount of information, including the obligatory egocentric (distance and direction) and allocentric (relationships among the objects) information, for object search. On the other hand, participants have varying perceptions of the optional information (e.g.  color and alert information).
@@ 5.1. RQ1: To what extent does ObjectFinder deliver the necessary information for effective object search?
Regarding the amount of information provided, three participants found it adequate for their needs. The average system feedback for route planning (F3) and scene description (F4) contains $62.90$ words, with a standard deviation of $19.11$. System feedback for open questions (F5) is relatively shorter. However, when asked about their preference for more or less information, responses were divided: half preferred more, while the other half favored less. We noted that preferences for information quantity relate to individual processing styles: while some participants could ignore excess information, others felt overwhelmed. The participants chose to receive more information, generally expressing a preference for as much as possible. As one participant noted, “as much as you can get. It’s completely blank for me, so the more I have, the better.” (P5). P4 expressed that “things that don’t interest me, I can just ignore.” Conversely, P6 and P7 explained their preference for less information, attributing it to not being accustomed to processing such a large amount of visual information and feeling overwhelmed by it. P8 suggested that the information provided could perhaps be reduced after getting the first overview: “it could maybe be a little bit less, so if you know once there are some objects on the table, then you don’t really need this information the second or third time unless you ask the system what’s on the table.” Additionally, we observed that the amount of information varies across functions. BeMyAI, used solely for scene description, received praise from three participants for its “detailed” information, although one deemed it excessive. In contrast, participants noted that our ObjectFinder delivered more information than necessary for efficient navigation.
@@
Participants (P6-8) highlighted the importance of localization information (distance and direction) after scene exploration. The interpretation of the distance should be clear and intuitive. P1 highlighted the usefulness of different distance measures noting, “I think the steps are good for like when it’s really near, so it’s just a few steps. But if there are longer distances. I think sometimes meters might be more useful […] So it sometimes has like $0.1$ meter […] you could have said ‘right in front of you’ or ‘one step’ like that.” Participants mentioned that the distances reported were inaccurate, appearing larger than they perceived. Upon reviewing the video, we observed that some discrepancies were caused by measurements taken from the head to the object, rather than horizontally. This inaccuracy may be more pronounced in our small indoor settings. To describe the direction, P5 suggested to “specify a $20$ centimeters margin”, noting that if an object is $10$ centimeters to the left or right, it’s still considered in the front.
Contextual information around the target is crucial for object search, such as relationships among surrounding objects and their location information. For example, P3 noted, “it’s good that the objects around it are announced, so you have some idea where the object you want to find is in relation to other objects.” A description should also include the distance, which says something about the user as a reference point to the object. P8 suggested to improve ObjectFinder by incorporating distance into the MLLM output, “ when you make the description, it holds everything, but it does not really talk about the distance.”
@@
5.2. RQ2: How do blind people perceive the detailed scene context that ObjectFinder provides to facilitate object search?
Participants were able to gain an overview of the scene, orient themselves, and explore search options within the detailed scene context provided by ObjectFinder.
There is some evidence that participants obtained more detailed environmental information through the use of ObjectFinder. P1 noted: “I think it was very good [to know] where I am. When it said where my target object was, it also told me the surrounding objects, because maybe if I find the surrounding object first, then I know, OK, I’m close. ” P6 noted the benefits of discovering other usable items (C2) and understanding their arrangement, “I didn’t know there was another chair there. It gives more information about objects and their arrangement.” Participants would like to get an overview of unfamiliar rooms without receiving too many details. P8 explained, “I could imagine if you really don’t know the room, and you want to first [have an] overview of what is in the room, it was pretty detailed. ”
@@
5.3. RQ3: How do blind individuals perceive ObjectFinder generally and in comparison to description- and detection-based systems?
Participants generally prefer ObjectFinder for object search over description- and detection-based systems because it enables users to specify targets actively and provides both egocentric and allocentric essential information, though some practice is required to become accustomed to it.
@@
Generally, five out of eight participants expressed excitement about ObjectFinder after exploring it. “I was very positively surprised at how good it works, and how easy it is to get information and to find myself in the scene.” (P1). ObjectFinder was also appreciated for accurate object searches without physical touch. Subsequently, participants rated the system’s helpfulness ($\mu=4.13,\sigma=0.641$) and their independence ($\mu=4.31,\sigma=1.01$) with ObjectFinder in an unfamiliar environment, with both ratings based on $5$-point Likert scales and averaged over $4$. Through ObjectFinder, participants discovered more than just target items, finding inspiration in unexpected objects, both small and large (C2). For example, “I’m sort of inspired because now I know what options I have, instead of walking in and just looking for my phone because I know it must be there, but I don’t know what else could be there […] I was inspired to look for, for example, bananas and stuff like that.”. (P1). However, P3 noted that searching with ObjectFinder might be slower than tactile exploration in our compact indoor setting, noted, “there are so many steps putting it up and searching for the thing, and maybe it doesn’t even find it (ObjectFinder_Base).” The search could be accelerated by using ObjectFinder alongside the cane, taking advantage of the cane’s large radius (P2).
@@
Regarding the differences between the commercial applications and the prototypes, participants valued that they could actively define the target objects with ObjectFinder, making it more reliable in searching for specific items. As P6 mentioned, “what I also liked about this system is speaking to it, I find that more targeted.”. P3 echoed this sentiment, highlighting the biggest advantage: “The biggest advantage […] contrary to BeMyAI is that you can specify what you want to find.” (C1). ”Furthermore, as previously mentioned, ObjectFinder delivers essential information for object searches that cannot be provided by either description-based or detection-based systems. This was more difficult using the existing commercial systems. For example, after using BeMyAI, P8 mentioned, ”BeMyAI had no information regarding the distances, so you have […] a less sense of location ”. As previously analyzed, egocentric localization information (distance and direction) and the allocentric relationships among surrounding objects are crucial for effective object search. The lack of localization information (P2, P8) and the absence of context regarding object relationships (P2) were mentioned as reasons why participants did not prefer Lookout, highlighting the potential for solutions that can combine multiple aspects of object search.
As shown in Table 4, the study also examined cognitive load, showing that it was relatively low across system types and scenarios. Retrieving keyframes for MLLM information and participant feedback reveals that the landscape orientation of the camera on the glasses, contrasting with participants’ usual practice of taking portrait photos using commercial applications, and the presence of relatively low furniture in the living room often result in the prototype failing to recognize the coffee table as an obstacle to the sofa. The exploratory experience with the prototype was not comprehensive enough. P7 mentioned that F4 and F5 are insufficiently tested in our study, and added: “one needs to get used to it, it requires some practice, and then the search will be very precise.” P5 mentioned that she was not accustomed to the earcon indicating to stand still.
@@
5.4. RQ4: What requirements do blind users consider important for an object search system, as their experiences with ObjectFinder show?
Participants basically suggested that efficiency is crucial in terms of interaction features and hardware for object search. Additionally, they considered object detection, localization, and route planning to be important functions of object search.
@@
Regarding target specification, P8 described the voice commands as “pretty, pretty easy”, while P4 expressed concerns with programs that solely accept voice input, suggesting “maybe it’s better to have the ability to switch between a list of maybe recognized objects and voice commands.” Besides the current earcons, P6 suggested creating an additional earcon for option confirmation, rather than using spoken text.
@@
We asked participants to rate the functions before and after using ObjectFinder and ObjectFinder_Base in the explored scenarios, in terms of their perceived importance and level of interest, as captured by Likert scores (Figure 10). From the ratings, the functions object detection, localization information, and route planning are important for the object search task ($\mu{>}4.0$), accompanied by a reduced standard error. Among the five functions evaluated, localization information was rated as the most important.
Compared with the commercial applications, P8 mentioned the advantage of ObjectFinder is “the distance and also the guiding function which is not really available for the other both apps.” (P8). We note that the software should be fast and make fewer mistakes. P2 mentioned that receiving explanations from a sighted person on BeMyEyes “works better because it’s without delay.” Half of the participants mentioned that frequent misidentifications of Lookout were not preferred. As P5 noted, “it misidentified objects […], like there’s no dishwasher,” and at times, the teapot was categorized as a helmet or mouse.